You cause a particle to move from point A, where the electric potential is 12.7 V, to point B, where the electric potential is − 29.7 V. Calculate the change that occurs in the particle's electrostatic potential energy, when the particle is an electron, a proton, a neutral hydrogen atom, and a singly ionized helium atom (i.e., lacking one electron from its neutral state).

Electron =

Proton =

Neutral Hydrogen Atom =

Electron =

Singly ionized helium atom (i.e., lacking one electron from its neutral state) =
Answer #1

va=12.7v

vb=-29.7v

change in potential=vb-va

= -29.7-12.7

=-42.4v

a)q=-1.6*10^-19c

electrostatic potential energy=potential difference*charge

=-1.6*10^-19*(-42.4)J

=67.8*10^-19J

b) for proton change in electrostatic potential energy is

-67.8*10^-19J

c)for neutral hydrogen atom change in electrostatic potential energy is zero (q=0c)

d)singly ionized helium atom q=1.6*10^-19c

change in electrostatic potential energy is =-67.8*10^-19J